ali can do a job in 9 hr, while susan can do the same job in 5 hr. how long would it take them to do the job if they worked together?
Answer by TimothyLamb(4379)

x = ali work rate (jobs/hr)
y = susan work rate (jobs/hr)
---
rate = work/time
time = work/rate
---
rate working together:
x + y = 1/9 + 1/5
---
time working together:
time = 1/(x + y) = 1/( 1/9 + 1/5 )
---
answer:
time = 3.21428571429 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php